From a piece of wire 30 cm long, it is required to bend a rectangle of the largest area. What are the dimensions of this rectangle


1
Expert's answer
2021-10-18T07:48:39-0400

Let "x=" the width of the rectangle.

Then the length of the rectangle will be "\dfrac{30-2x}{2}=15-x."

The area of the rectangle is


"A=A(x)=x(15-x), 0<x<15"

"A(x)=15x-x^2=-(x^2-15x+56.25)+56.25"

"=-(x-7.5)^2 +56.25"

Vertex "(7.5, 56.25)."

The quadratic function "A(x)" has the absolute maximum with value of "56.25" at "x=7.5."


width "7.5" cm


length "7.5" cm